Twelve less than one fourth of the year I was born is equal to nine times my age plus forty eight.
Answer by jorel555(1290)

Assuming we're using your age this year, 2016, and letting n be your birth year; then:
(n/4)-12=9(2016-n)+48
(n/4)-12=18144-9n+48
9n+(n/4)=18204
36n+n=72186
n=1968!!!!!!!!
If you were born in 1968, then you are 2016-1968, or 48 years old now!!!!!!!!!!!